\begin{thebibliography}{99}

\bibitem{Golan1999}
J.~S. Golan,
\emph{Semirings and Their Applications},
Springer, Dordrecht, 1999.

\bibitem{HebischWeinert1998}
U.~Hebisch and H.~J. Weinert,
\emph{Semirings: Algebraic Theory and Applications in Computer Science},
World Scientific, Singapore, 1998.

\bibitem{Bourne1951}
S.~R. Bourne,
``The Theory of Semirings with Applications in Mathematics,''
\emph{Proceedings of the London Mathematical Society},
vol.~3, no.~1, pp.~193--212, 1951.

\bibitem{Nobusawa1964}
N.~Nobusawa,
``On Ternary Semigroups,''
\emph{Mathematical Journal of Okayama University},
vol.~13, pp.~1--7, 1964.

\bibitem{Post1940}
E.~Post,
``Polyadic Groups,''
\emph{Transactions of the American Mathematical Society},
vol.~48, no.~2, pp.~208--350, 1940.

\bibitem{Dornte1928}
W.~D\"ornte,
``Untersuchungen \"uber einen verallgemeinerten Gruppenbegriff,''
\emph{Mathematische Zeitschrift},
vol.~29, pp.~1--19, 1928.

\bibitem{Hosszu1963}
M.~Hossz\'u,
``On n-ary Groups,''
\emph{Acta Mathematica Academiae Scientiarum Hungaricae},
vol.~14, pp.~239--246, 1963.

\bibitem{Balaban1992}
A.~T. Balaban (ed.),
\emph{Chemical Graph Theory},
CRC Press, Boca Raton, 1992.

\bibitem{Trinajstic1992}
N.~Trinajsti\'c,
\emph{Chemical Graph Theory},
CRC Press, Boca Raton, 1992.

\bibitem{GutmanPolansky1986}
I.~Gutman and O.~E. Polansky,
\emph{Mathematical Concepts in Organic Chemistry},
Springer, Berlin, 1986.

\bibitem{BanerjeeLavine1994}
A.~Banerjee and B.~Lavine,
``A Mathematical Approach to Chemical Reactivity,''
\emph{Journal of Mathematical Chemistry},
vol.~15, pp.~235--250, 1994.

\bibitem{King2001}
R.~B. King,
``Applications of Graph Theory and Topology in Inorganic Cluster and Coordination Chemistry,''
\emph{Chemical Reviews},
vol.~101, no.~11, pp.~2873--2920, 2001.

\bibitem{Tropashko2006}
V.~Tropashko,
``Algebraic Extensions of Ternary Relations: A Semiring Framework,''
\emph{Information Processing Letters},
vol.~98, no.~3, pp.~110--115, 2006.

\end{thebibliography}

\documentclass[12pt,a4paper]{article}

\usepackage[T1]{fontenc}
\usepackage[utf8]{inputenc}
\usepackage{amsmath,amssymb,amsfonts,amsthm}
\usepackage{mathtools}
\usepackage{hyperref}
\usepackage{enumitem}
\usepackage{geometry}
\newtheorem{definition}{Definition}[section]

\newtheorem{proposition}[definition]{Proposition}

\begin{document}

\title{\textbf{Axiomatic Foundations of Chemical Systems as Ternary $\Gamma$-Semirings}}

\author{\small
Chandrasekhar Gokavarapu$^{1,2}$\hspace{5mm} Venkata Rao Kaviti$^{3}$ \\
\small  Srinivasa Rao Thirunagari $^{4}$ \hspace{5mm} Dr D Madhusudhana Rao$^{5,6}$\\[2pt]
 \small $^1$Lecturer in Mathematics,
Government College (A), Rajahmundry, A.P., India\\
 \small $^2$Research Scholar, Department of Mathematics,
Acharya Nagarjuna University, Guntur, A.P., India\\
\texttt{\small chandrasekhargokavarapu@gmail.com}\\ 
 \small $^3$Lecturer in Chemistry,
Government College (A), Rajahmundry, A.P., India\\
 \small $^4$Lecturer in Chemistry,
Government College (A), Rajahmundry, A.P., India\\
\small $^5$Lecturer in  Mathematics, Government College For Women(A), Guntur, Andhra Pradesh, India,\\
\small $^6$Research Supervisor, Dept.  of Mathematics,Acharya Nagarjuna University, Guntur, A.P., India,\\
\texttt{\small dmrmaths@gmail.com}}
\date{}
\maketitle
\begin{abstract}
\begin{sloppypar}
Chemical transformations depend not only on the identities of the 
reacting species but also on the catalytic, environmental, and 
intermediate conditions under which they occur. Classical binary 
reaction formalisms usually treat such conditions as external 
annotations, which obscures the genuinely multi-state and 
multi-parameter character of real chemical processes.

In this paper we introduce an axiomatic framework in which a chemical 
system is modelled by a ternary $\Gamma$-semiring. The elements of the 
state set represent chemical states, while the parameter set encodes 
catalytic and environmental conditions. A $\Gamma$-dependent ternary 
operation is used to describe mediated transformations, treating 
reactants, intermediates, and mediators as intrinsic arguments of the 
transformation law.

We develop the algebraic axioms governing these mediated interactions 
and interpret their associativity, distributivity, and 
$\Gamma$-linearity in terms of multi-step pathways, parallel processes, 
and controlled environmental dependence. We introduce chemical ideals 
and $\Gamma$-ideals as algebraic structures modelling reaction-closed 
subsystems and pathway-stable domains, and study their prime and 
semiprime forms. Homomorphisms between TGS-chemical systems are shown to 
preserve reaction pathways and describe consistent changes of chemical 
environment.

Abstract examples from catalysis, thermodynamic phase control, and 
field-induced quantum transitions illustrate how familiar chemical 
phenomena fit within this framework. The resulting theory provides a 
unified algebraic foundation for multi-parameter chemical behaviour and 
establishes the structural basis for subsequent developments involving 
kinetics, geometric methods, and computational or AI-assisted models.
\end{sloppypar}
\end{abstract}

\tableofcontents

\section{Introduction}

Chemical systems have long served as a rich source of intuition and
examples for mathematics, while mathematical structures have, in turn,
provided increasingly refined languages for describing reactivity,
stability, and transformation in chemistry.
Classical formalisms in chemical kinetics, thermodynamics, and quantum
chemistry typically encode reactions in terms of \emph{binary}
combinations of species (see \cite{BanerjeeLavine1994} for mathematical models of reactivity)
,
\[
  A + B \longrightarrow C,
\]
with additional information---such as catalysts, solvents, temperature,
pressure, or external fields---being attached as annotations to the
reaction arrow rather than as intrinsic components of the algebraic
operation itself.
This viewpoint is extremely successful in many settings, but it obscures
the genuinely multi-parameter and multi-state nature of real chemical
transformations.

In practice, a reaction pathway is rarely determined solely by the
identities of the reacting species.
Instead, it is governed by a constellation of mediating factors:
catalysts that open or close pathways, solvent environments that
stabilize intermediates, pressure--temperature conditions that reshape
energy landscapes, and external fields that deform quantum states.
These ingredients do not simply modify a pre-existing binary operation;
they participate structurally in how chemical states are transformed.
From an algebraic viewpoint, this suggests that the primitive operation
underlying chemical change should be a higher-arity map (higher-arity algebraic structures were earlier studied in 
\cite{Dornte1928,Post1940,Nobusawa1964,Hosszu1963})
 that treats
mediators on the same footing as the states they control (compare with classical semiring frameworks \cite{Golan1999,HebischWeinert1998})
.

The aim of the present paper is to make this intuition precise.
We propose an axiomatic framework in which a chemical system is modeled
by a ternary~$\Gamma$-semiring, and in which the fundamental
reaction-like transformation is encoded by a $\Gamma$-dependent ternary
operation
\[
  [A,\alpha,B,\beta,C] \in S,
\]
where $A,B,C \in S$ represent chemical states and
$\alpha,\beta \in \Gamma$ represent mediating parameters.
This construction elevates catalysts, solvents, and environmental
conditions from external labels to algebraically active inputs, thereby
providing a unified structure in which multi-state, multi-parameter(for general algebraic perspectives, see \cite{Bourne1951})

interactions can be studied with the full precision of modern algebra.

\subsection{Motivation}

The starting point for our work is the observation that classical
reaction notation is intrinsically binary both in syntax and in its
implicit algebraic interpretation.
A formal reaction of the form
\[
  A + B \longrightarrow C
\]
suggests an underlying binary operation that takes a pair of input
states $(A,B)$ and produces an output state~$C$.
When catalysts or conditions are present, one typically writes
\[
  A + B 
  \xrightarrow[\text{solvent}]{\text{catalyst},\,T,p} 
  C,
\]
but the additional data are carried outside the core operation;
they do not enter as arguments of the algebraic map itself.
In particular, the difference between a catalyzed and an uncatalyzed
reaction, or between two reactions run under distinct temperature
profiles, is not reflected at the level of the algebraic arity.

However, empirical chemistry shows that these ``external'' features are
often decisive.
The presence or absence of a catalyst can completely alter both the
available pathways and the final products.
Solvents stabilize different intermediates, reshaping the energy
landscape.
Pressure and temperature selectively favor certain phases or reaction
channels, while electric or magnetic fields modify quantum states and
transition probabilities.
All of these effects are not accidental decorations but intrinsic
components of how chemical states transform.

From a structural point of view, this suggests that chemical systems
should be viewed as \emph{mediated} transformation systems:
the outcome of an interaction between states is mediated by additional
parameters that influence, constrain, or enable certain transitions.
Instead of encoding this mediation by enlarging the state space in an
ad hoc manner, it is natural to treat the mediators as elements of a
separate set~$\Gamma$ and to allow the basic operation to depend
explicitly on them.
A higher-arity algebraic system whose fundamental operation
\[
  [A,\alpha,B,\beta,C]
\]
takes both states and mediators as arguments is then a natural candidate
for formalizing chemical behaviour.

This shift in perspective has several conceptual advantages.
First, it allows us to distinguish chemically between different uses of
the same species under distinct conditions without artificially
duplicating the state space.
Second, it creates a direct route for encoding multi-step and
cooperative phenomena: ternary operations can be iterated and composed
in a way that keeps track of how mediators combine or interact.
Third, it aligns chemical reasoning with modern algebraic practices,
where higher-arity operations and parameterized structures play a
central role in understanding complex systems.

\subsection{Why Ternary \texorpdfstring{$\Gamma$}-Semirings?}

The abstract notion of a ternary~$\Gamma$-semiring provides a
particularly suitable environment for realizing the above programme.
At a formal level, a ternary~$\Gamma$-semiring consists of a set~$S$,
a parameter set~$\Gamma$, and a $\Gamma$-dependent ternary operation
\[
  [\,\cdot\,,\cdot\,,\,\cdot\,,\cdot\,,\,\cdot\,]
  \colon S \times \Gamma \times S \times \Gamma \times S \longrightarrow S
\]
that satisfies appropriate associativity, distributivity, and
$\Gamma$-linearity conditions.
When $S$ is interpreted as a space of chemical states and $\Gamma$ as a
space of mediators (such as catalysts, solvents, or thermodynamic
controls), the value
\[
  [A,\alpha,B,\beta,C]
\]
can be read as the resulting state of a mediated transformation in which
$A,B,C$ interact under the influence of $\alpha$ and~$\beta$.

Several features of ternary~$\Gamma$-semirings align naturally with
chemical behaviour:

\begin{itemize}
  \item \emph{Catalyst-dependent reactions.}
  Mediators in $\Gamma$ can represent catalysts, so that different
  catalytic scenarios correspond to different choices of $\alpha$ and
  $\beta$, even when the underlying species $A,B,C$ are fixed.

  \item \emph{Solvent and environment effects.}
  Solvents and bulk environmental parameters can be encoded as elements
  of~$\Gamma$, allowing changes in solvent or ambient medium to be
  reflected directly as changes in the operators governing~$S$.

  \item \emph{Pressure and temperature conditions.}
  Thermodynamic variables can be grouped into mediating parameters,
  making it possible to distinguish transformations that are identical
  in stoichiometry but distinct in their pressure--temperature regimes.

  \item \emph{Multi-species interactions.}
  The ternary operation simultaneously involves three states of~$S$,
  permitting the modelling of complex elementary steps, cooperative
  effects, or intermediate formation within a single algebraic act.
\end{itemize}

From an algebraic standpoint, the $\Gamma$-dependence provides a
controlled way to encode families of reaction laws indexed by
conditions, while the ternary nature reflects the intrinsically
multi-input character of mediated transformations.
The ternary~$\Gamma$-semiring therefore emerges as a natural and
flexible candidate for an axiomatic definition of chemical systems.

\subsection{Contribution of this paper}

In this work we develop a systematic axiomatic theory of chemical
systems based on ternary~$\Gamma$-semirings.
More precisely, we proceed along the following lines:

\begin{itemize}
  \item We introduce the notion of a \emph{TGS-chemical system},
  defined as a ternary~$\Gamma$-semiring whose elements are interpreted
  as chemical states and whose $\Gamma$-indexed ternary operation
  encodes mediated transformations of those states.
  The central object of study is the map
  \[
    [A,\alpha,B,\beta,C] \in S,
  \]
  which we interpret as the resulting state of a ternary interaction
  between $A,B,C \in S$ under mediators $\alpha,\beta \in \Gamma$.

  \item We formalize the reaction operation as a $\Gamma$-mediated
  ternary map and specify axioms that capture associativity,
  distributivity, and compatibility with the $\Gamma$-structure in a
  way that reflects multi-step reactions, parallel pathways, and
  composite environments.

  \item We develop the structural theory of TGS-chemical systems,
  introducing and analyzing suitable notions of ideals and
  $\Gamma$-ideals that correspond to chemically meaningful subsystems
  and reaction-closed families of states.
  In particular, we study prime and semiprime ideals in this setting
  and interpret them in terms of irreducible or stability properties of
  reaction networks.

  \item We investigate homomorphisms of TGS-chemical systems as
  structure-preserving maps between chemical environments.
  These homomorphisms provide a natural language for comparing and
  transporting reaction laws between different systems, and for
  formalizing operations such as changing solvent, adjusting
  environmental conditions, or embedding a subsystem into a larger
  chemical context.

  \item Throughout, we illustrate the theory with examples that show
  how catalyzed reactions, solvent effects, phase transitions, and other
  chemically relevant phenomena can be encoded within the TGS framework,
  thereby demonstrating that the proposed axioms are not merely formal
  but admit a concrete interpretation in chemical practice.
\end{itemize}

Taken together, these contributions establish a unified algebraic picture
in which chemical states, mediators, and transformations are treated
within a single ternary~$\Gamma$-semiring structure.
This provides a foundation on which further developments---including
kinetic refinements, computational models, and connections to symbolic
reasoning and machine-assisted chemistry---can be built in subsequent
work.

\section{Preliminaries on Ternary \texorpdfstring{$\Gamma$}-Semirings}

In this section we recall the algebraic notions needed throughout the
paper.  Our treatment follows standard practice in the theory of
higher-arity algebraic systems, adapted to the setting of
$\Gamma$-parameterized ternary operations.  No chemical interpretation
is introduced here; the objective is purely structural.  All subsequent
sections will build on these foundations.

\subsection{Ternary operations}

A \emph{ternary operation} on a set $S$ is a map
\[
  \mu \colon S \times S \times S \longrightarrow S,
\](see \cite{Dornte1928,Post1940}).
which assigns to each triple $(A,B,C)$ an element $\mu(A,B,C)\in S$.
Ternary operations generalize the familiar notion of binary operations
by allowing three inputs to participate simultaneously in the formation
of a new element.  In the presence of additional structure, such as a
parameter set or distributive laws, ternary operations serve as the
basic building blocks for higher-arity semigroup or semiring-like
systems.

\subsection{\texorpdfstring{$\Gamma$}-sets and parameterized operations}

Throughout this paper, $\Gamma$ denotes a nonempty set whose elements
act as \emph{mediating parameters}.
A \emph{$\Gamma$-set} is simply a pair $(S,\Gamma)$ consisting of a set
$S$ together with an external parameter set~$\Gamma$.
The elements of $\Gamma$ do not act on $S$ directly unless a specific
operation is specified; instead, they serve as indices governing how
elements of $S$ combine.

In particular, a \emph{$\Gamma$-parameterized ternary operation} on~$S$
is a map
\[
  [\,\cdot\,,\cdot\,,\,\cdot\,,\cdot\,,\,\cdot\,]
  \colon S \times \Gamma \times S \times \Gamma \times S
  \longrightarrow S,
\]
where the parameters in~$\Gamma$ may influence the resulting element in
a nontrivial way.
This form of parameterization is essential for modelling situations in
which the behaviour of a ternary interaction depends on contextual or
environmental data.

\subsection{Ternary \texorpdfstring{$\Gamma$}-semirings}

We now introduce the central notion used in this work \cite{Golan1999,HebischWeinert1998}
.

\begin{definition}
A \emph{ternary $\Gamma$-semiring} is a triple
$(S,\Gamma,[\,])$ consisting of a nonempty set~$S$, a nonempty parameter
set~$\Gamma$, and a $\Gamma$-parameterized ternary operation
\[
  [\, , \, , \, , \, , \,]
  \colon S \times \Gamma \times S \times \Gamma \times S \to S,
\]
satisfying the following axioms for all
$A,B,C,D,E \in S$ and all $\alpha,\beta,\gamma,\delta \in \Gamma$:
\begin{enumerate}[label=\textup{(T\arabic*)}]
  \item \textbf{Associativity.}
  The operation is associative in the sense that
  \[
    [A,\alpha,[B,\beta,C,\gamma,D],\delta,E]
    =
    [[A,\alpha,B,\beta,C],\gamma,D,\delta,E],
  \]
  whenever the expressions are formed.
  This ensures that iterated ternary combinations are well defined.

  \item \textbf{$\Gamma$-linearity.}
  For fixed internal arguments, the dependence on the parameters
  $\alpha,\beta$ is compatible with the $\Gamma$-structure.
  (The specific linearity or compatibility conditions imposed on
  $\Gamma$ will be detailed when required for structural results.)

  \item \textbf{Distributivity.}
  The ternary operation distributes over itself in each argument in the
  appropriate higher-arity analogue of semiring distributivity.
  For instance,
  \[
    [A,\alpha,B,\beta,[C,\gamma,D,\delta,E]]
    =
    [[A,\alpha,B,\beta,C],\gamma,D,\delta,E]
    =
    [A,\alpha,[B,\beta,C,\gamma,D],\delta,E],
  \]
  with analogous conditions holding in the remaining positions.
  These distributivity relations guarantee that the operation behaves
  coherently when nested.
\end{enumerate}
\end{definition}

The axioms above give a flexible framework in which ternary interactions
can be iterated, nested, and composed while respecting a fixed set of
mediating parameters.(related ternary operations appear in \cite{Nobusawa1964})

In later sections we will interpret $S$ as a set of chemical states and
$\Gamma$ as a set of mediators (such as catalysts, solvents, or
thermodynamic conditions), with the ternary operation modelling
parameter-dependent transformations.
At this stage, however, we treat $(S,\Gamma,[\,])$ as a purely algebraic
object, postponing chemical meaning until the core definitions of
TGS-chemical systems are introduced.

\section{Chemical Systems as Ternary \texorpdfstring{$\Gamma$}-Semirings}

We now introduce the central conceptual framework of the paper: a 
chemical system viewed as a ternary~$\Gamma$-semiring whose elements 
represent chemical states and whose mediators encode the environmental 
or catalytic factors influencing their transformations.
While the preceding section provided the purely algebraic foundation,
our goal here is to explain how these structures naturally model 
multi-parameter, multi-state chemical behaviour.

\subsection{Chemical interpretation of \texorpdfstring{$S$}{S} and \texorpdfstring{$\Gamma$}{Gamma}}

Let $S$ be a nonempty set.
In the context of chemical systems, we interpret the elements of~$S$ as 
\emph{chemical states}.
The notion of a state is intentionally broad and may encompass:

\begin{itemize}
  \item molecular configurations or species identities;
  \item concentration levels in a reaction mixture;
  \item phase descriptors (solid, liquid, gas, plasma);
  \item electronic, vibrational, or quantum mechanical states;
  \item intermediate structures arising during reaction pathways.
\end{itemize}

Thus, $S$ serves as the universe within which chemically meaningful
objects reside.This perspective aligns with classical mathematical chemistry 
frameworks that treat chemical structure and states using abstract 
mathematical representations (see \cite{Balaban1992,Trinajstic1992,GutmanPolansky1986}).

Let $\Gamma$ be a nonempty parameter set.
Its elements are interpreted as \emph{mediators}, representing 
conditions or influences under which chemical interactions occur.  
Typical examples include:

\begin{itemize}
  \item catalysts and co-catalysts;
  \item solvent environments;
  \item pressure and temperature conditions;
  \item electromagnetic or external field parameters;
  \item pH, ionic strength, or other environmental controls.
\end{itemize}

These mediators do not transform chemical states directly; rather, they
govern or modulate the transformation rules encoded by the ternary
operation defined below.  
In this sense, $(S,\Gamma)$ forms the structural substrate of a 
chemical system.Such a parametrization of environmental and catalytic conditions 
is consistent with algebraic treatments of ternary and mediated 
transformations in other settings (compare \cite{Tropashko2006}).

\subsection{Core chemical operation}

The essential ingredient of a TGS-chemical system is a  
$\Gamma$-parameterized ternary operation
\[
  [\,\cdot\,,\cdot\,,\,\cdot\,,\cdot\,,\,\cdot\,]
  \colon S \times \Gamma \times S \times \Gamma \times S
  \longrightarrow S,
\]
which assigns, to each triple of states $A,B,C \in S$ and each pair of 
mediators $\alpha,\beta \in \Gamma$, a resulting state $D \in S$.
We write this compactly as
\[
  [A,\alpha,B,\beta,C] = D.
\]

Chemically, this is interpreted as follows:
\begin{itemize}
  \item $A$ is an initial or reactant state;
  \item $B$ is an interacting state, possibly another reactant or an 
        intermediate;
  \item $C$ is a subsequent state, often representing an intermediate or 
        transition configuration;
  \item $\alpha,\beta$ encode mediating conditions (catalysts, solvents, \\ temperature/pressure regimes, or external fields);
  \item $D$ is the resulting state after the mediated interaction of 
        $A$, $B$, and $C$ under parameters $\alpha$ and $\beta$.
\end{itemize}

This notation may be viewed as a symbolic representation of a 
parameter-dependent reaction pathway:
\[
  A \xrightarrow{\alpha} B \xrightarrow{\beta} C \quad\rightsquigarrow\quad D,
\]
in which the overall transformation is encoded by the ternary 
$\Gamma$-operation.
Unlike the classical binary reaction form $A+B \to C$, this framework 
treats mediators as intrinsic arguments of the operation rather than 
external labels.This sharply contrasts with binary mathematical models of reactivity 
commonly used in algebraic treatments of chemical transformations 
(see \cite{BanerjeeLavine1994}).

This allows chemically distinct processes that share the same 
stoichiometry but differ in conditions to be represented distinctly at 
the algebraic level.

\subsection{Axioms for chemical TGS}

The axioms of a ternary~$\Gamma$-semiring, introduced earlier in a 
purely algebraic setting, acquire a natural chemical interpretation 
when applied to the present framework.
We summarize the interpretative content of the main axioms below.

\paragraph*{(1) Associativity and multi-step reactions.}
The associativity axiom ensures that the outcome of a sequence of 
mediated transformations is independent of the order in which the 
ternary combinations are grouped.
Chemically, this corresponds to the fact that a multi-step reaction 
pathway
\[
  A \longrightarrow B \longrightarrow C \longrightarrow D \longrightarrow E
\]
admits a coherent overall description, regardless of whether one 
groups intermediate steps as $(A\to B\to C)$ followed by $(C\to D\to E)$ 
or uses another valid decomposition.
Thus, associativity provides an algebraic representation of 
multi-step or multi-intermediate reaction processes.

\paragraph*{(2) $\Gamma$-linearity and scaling of conditions.}
The $\Gamma$-linearity condition expresses compatibility between the 
mediators and the ternary operation.
While no specific algebraic structure on $\Gamma$ is imposed at this 
stage, the general principle is that variations or combinations of 
catalytic or environmental parameters correspond to predictable or 
structured variations in the resulting state.
From a chemical standpoint, increasing catalyst concentration, changing 
solvent polarity, or adjusting temperature should influence reaction 
behaviour in a manner consistent with the dependence encoded by the 
operation $[A,\alpha,B,\beta,C]$.

\paragraph*{(3) Distributivity and parallel reactions.}
The distributivity axioms capture the idea that the ternary $\Gamma$-operation 
behaves coherently when nested or combined with itself.
Chemically, this reflects the presence of parallel or branching reaction 
pathways.
For example, if $C$ can arise from multiple competing intermediates 
or if the environment induces branching in the transformation sequence, 
the distributive laws ensure that such behaviour is represented in a 
controlled algebraic manner.
Distributivity therefore encodes the superposition or recombination of 
reaction channels.

\vspace{2mm}

Together, these axioms allow ternary~$\Gamma$-semirings to model 
chemical systems in which states evolve under the influence of 
environmental conditions, catalysts, and other mediating factors.
The remainder of the paper develops the structural theory of such 
systems and illustrates how classical and nonclassical chemical processes 
fit naturally within the TGS framework.

\section{Structural Theory of TGS-Chemical Systems}

In this section we develop the basic structural theory of 
TGS-chemical systems.
Our aim is to identify those subsets of the state space~$S$ that behave
as chemically meaningful subsystems, closed under reaction and stable 
under the mediating parameters~$\Gamma$.
These subsets will be formalized as various kinds of ideals, and their 
properties will be interpreted in terms of reaction networks and 
pathways.

Throughout, $(S,\Gamma,[\,])$ denotes a fixed ternary~$\Gamma$-semiring
equipped with the chemical interpretation of Section~3.

\subsection{Chemical ideals}

We first single out subsets that are internally closed under the 
reaction operation and, in a stronger form, absorb interactions with 
the ambient system in a controlled way.

\begin{definition}[Reaction-closed subset]
A nonempty subset $R \subseteq S$ is called \emph{reaction-closed}
if for all $A,B,C \in R$ and all $\alpha,\beta \in \Gamma$,
\[
  [A,\alpha,B,\beta,C] \in R.
\]
\end{definition}
This notion generalizes closure concepts appearing in classical semiring
structures (compare \cite{Bourne1951,Golan1999}), but adapted to the
mediated ternary operation governing chemical interactions.

In chemical terms, a reaction-closed subset represents a collection of
states that, once present together in the system, can only yield states
that remain within the same collection, irrespective of the mediating
conditions.
Such a subset may be viewed as a self-contained reaction universe:
all internally accessible states via the ternary operation stay inside $R$.

Reaction-closedness captures purely internal behaviour.
To model interaction with the surrounding system, we require an
absorption property.

\begin{definition}[Chemical ideal]
A nonempty subset $I \subseteq S$ is called a \emph{chemical ideal}
if it satisfies the following conditions:
\begin{enumerate}[label=\textup{(C\arabic*)}]
  \item \emph{Internal closure:} for all $A,B,C \in I$ and all
        $\alpha,\beta \in \Gamma$,
        \[
          [A,\alpha,B,\beta,C] \in I;
        \]
  \item \emph{Boundary absorption:} for all $A,C \in I$, all $B \in S$
        and all $\alpha,\beta \in \Gamma$,
        \[
          [A,\alpha,B,\beta,C] \in I.
        \]
\end{enumerate}
\end{definition}
The internal closure and absorption properties reflect the role of
one- and two-sided ideals in semiring theory 
(\cite{Bourne1951,Golan1999}), extended here to the ternary $\Gamma$-interaction
and its chemical interpretation.

Condition~(C1) states that $I$ is reaction-closed in the sense defined
above.
Condition~(C2) expresses that if a mediated transformation begins and
ends in $I$, then the entire effect of any intervening state $B$ and 
any mediating parameters $\alpha,\beta$ remains confined to~$I$.
Chemically, $I$ can be thought of as a subsystem that is closed under
all internal reactions and stable under any process that connects two
of its states, even when intermediate species from outside $I$ are 
involved.

\begin{proposition}
The intersection of any family of chemical ideals in $S$ is again a 
chemical ideal.
\end{proposition}

\begin{proof}
Let $\{I_j\}_{j \in J}$ be a family of chemical ideals and set
$I := \bigcap_{j \in J} I_j$.
Since each $I_j$ is nonempty, the intersection is either empty or
nonempty; if empty, it is excluded from consideration, so we assume
$I \neq \varnothing$.
Let $A,B,C \in I$ and $\alpha,\beta \in \Gamma$.
Then $A,B,C \in I_j$ for every $j$, and by (C1) in each $I_j$ we have
$[A,\alpha,B,\beta,C] \in I_j$ for all $j$.
Hence $[A,\alpha,B,\beta,C] \in \bigcap_{j} I_j = I$, so (C1) holds
for~$I$.

Similarly, let $A,C \in I$, $B \in S$ and $\alpha,\beta \in \Gamma$.
Then $A,C \in I_j$ for every $j$, and by (C2) in each $I_j$ we obtain
$[A,\alpha,B,\beta,C] \in I_j$ for all $j$.
Therefore $[A,\alpha,B,\beta,C] \in I$, and (C2) holds.
Thus $I$ is a chemical ideal.
\end{proof}

This result shows that chemical ideals form a complete lattice under
intersection, providing a natural hierarchy of chemically stable
subsystems inside a given TGS-chemical system.Such hierarchical decompositions reflect analogous structural 
decomposition phenomena in classical semiring theory 
(see \cite{HebischWeinert1998}).

\subsection{$\Gamma$-ideals and reaction pathways}

The previous notion focuses on subsets of $S$ that are stable under
interactions involving their boundary states.
We now refine this by distinguishing the role of a single participating
state and allowing the other states to range freely.
This leads to a $\Gamma$-ideal structure, reflecting how certain states
control or channel reaction pathways.

\begin{definition}[$\Gamma$-ideals]
A nonempty subset $J \subseteq S$ is called:
\begin{enumerate}[label=\textup{(G\arabic*)}]
  \item a \emph{left $\Gamma$-ideal} if for all $X \in J$, all
        $A,B \in S$ and all $\alpha,\beta \in \Gamma$,
        \[
          [X,\alpha,A,\beta,B] \in J;
        \]
  \item a \emph{right $\Gamma$-ideal} if for all $X \in J$, all
        $A,B \in S$ and all $\alpha,\beta \in \Gamma$,
        \[
          [A,\alpha,B,\beta,X] \in J;
        \]
  \item a \emph{middle $\Gamma$-ideal} if for all $X \in J$, all
        $ A,C \in S $ and all $\alpha,\beta \in \Gamma$,
        \[
          [A,\alpha,X,\beta,C] \in J;
        \]
  \item a \emph{(two-sided) $\Gamma$-ideal} if it is simultaneously
        a left, right, and middle $\Gamma$-ideal.
\end{enumerate}
\end{definition}

Loosely speaking, a left $\Gamma$-ideal is stable under all transformations
in which one of its elements appears in the first argument position,
and similarly for right and middle $\Gamma$-ideals.
A two-sided $\Gamma$-ideal is stable under all ternary interactions in
which at least one position is occupied by an element of the ideal.

Chemically, these notions correspond to different forms of control over
reaction pathways:
\begin{itemize}
  \item a left $\Gamma$-ideal collects states that, once present as
        ``initiators'' of interactions, always lead back into the same
        collection, regardless of what they interact with;
  \item a right $\Gamma$-ideal behaves analogously for ``terminal''
        positions, capturing states that cannot be escaped once they
        appear as final products;
  \item a middle $\Gamma$-ideal models states that, when acting as
        intermediates, keep the system confined to a specific region of
        the state space;
  \item a two-sided $\Gamma$-ideal encodes a robustly closed set of
        states that controls and absorbs reaction pathways in all 
        three positions.
\end{itemize}

We now describe reaction pathways in this setting.

\begin{definition}[Reaction pathway]
Let $(S,\Gamma,[\,])$ be a TGS-chemical system.
A \emph{reaction pathway} of length $n \ge 1$ is a finite sequence
\[
  (X_0,X_1,\dots,X_n)
\]
of elements of $S$ such that for each $k = 1,\dots,n$ there exist
$A_k,B_k \in S$ and $\alpha_k,\beta_k \in \Gamma$ with
\[
  X_k = [A_k,\alpha_k,B_k,\beta_k,X_{k-1}]
  \quad\text{or}\quad
  X_k = [X_{k-1},\alpha_k,A_k,\beta_k,B_k]
  \quad\text{or}\quad
  X_k = [A_k,\alpha_k,X_{k-1},\beta_k,B_k].
\]
The element $X_0$ is the \emph{source} and $X_n$ the \emph{target} of
the pathway.
\end{definition}

This notion captures the idea that chemical evolution proceeds through
a chain of mediated ternary interactions, with each step determined by
a choice of two companion states and a pair of mediators.

\begin{proposition}
Let $J \subseteq S$ be a two-sided $\Gamma$-ideal.
If a reaction pathway $(X_0,\dots,X_n)$ satisfies $X_0 \in J$, then
$X_k \in J$ for all $k=0,\dots,n$.
\end{proposition}

\begin{proof}
We argue by induction on $k$.
For $k=0$ this is true by assumption.
Suppose $X_k \in J$ for some $0 \le k < n$.
By definition of a reaction pathway, $X_{k+1}$ is obtained from $X_k$
by one of the following forms:
\[
  X_{k+1} = [X_k,\alpha,A,\beta,B],\quad
  X_{k+1} = [A,\alpha,B,\beta,X_k],\quad
  X_{k+1} = [A,\alpha,X_k,\beta,B],
\]
for suitable $A,B \in S$ and $\alpha,\beta \in \Gamma$.
Since $J$ is a two-sided $\Gamma$-ideal, each of these expressions
belongs to $J$ whenever $X_k \in J$.
Hence $X_{k+1} \in J$, completing the induction.
\end{proof}

Chemically, this proposition states that once a system enters a
two-sided $\Gamma$-ideal, all states reachable via reaction pathways
remain confined within that ideal.
Thus, two-sided $\Gamma$-ideals model reaction basins or domains in
which the chemistry is dynamically trapped under the available
mediators.

\subsection{Prime and semiprime chemical ideals}

To understand how reaction activity distributes across the state space,
it is useful to introduce notions of primeness and semiprimeness that 
adapt classical ideal-theoretic concepts to the ternary 
$\Gamma$-setting.

\begin{definition}[Prime chemical ideal]
A proper chemical ideal $P \subsetneq S$ is called \emph{prime} if
whenever
\[
  [A,\alpha,B,\beta,C] \in P
\]
for some $A,B,C \in S$ and $\alpha,\beta \in \Gamma$, then at least one
of $A,B,C$ lies in $P$.
\end{definition}

This notion extends the classical understanding of prime ideals in 
semiring theory, where primeness forbids internal factorization of 
elements outside the ideal (compare \cite{Golan1999}).

Thus, a prime chemical ideal cannot contain the result of a mediated
interaction without ``detecting'' the presence of one of its
participants.
Chemically, $P$ behaves like a region of the state space whose boundary
is sufficiently sharp that it cannot be entered as the product of a
reaction between three states all lying outside~$P$.
In this sense, a prime chemical ideal captures a subsystem in which 
one interaction or family of interactions dominates access to its 
interior.

\begin{definition}[Semiprime chemical ideal]
A chemical ideal $I \subseteq S$ is called \emph{semiprime} if for every
$A \in S$ the following implication holds:
if
\[
  [A,\alpha,A,\beta,A] \in I
  \quad\text{for all }\alpha,\beta \in \Gamma,
\]
then $A \in I$.
\end{definition}
This condition generalizes the classical semiprime property in semiring 
theory, where self-combinations inside an ideal force membership of the 
element itself (see \cite{Golan1999}).

Here, $[A,\alpha,A,\beta,A]$ may be understood as a self-interaction or
self-combination of the state $A$ under all possible mediators.
The definition says that if every such self-interaction of $A$ falls 
inside $I$, then $A$ itself must already belong to~$I$.
Semiprimeness thus prevents the existence of ``hidden'' states outside 
$I$ whose entire mediated self-dynamics is trapped within~$I$.

\begin{proposition}
Every prime chemical ideal is semiprime.
\end{proposition}

\begin{proof}
Let $P$ be a prime chemical ideal and suppose that
$[A,\alpha,A,\beta,A] \in P$ for all $\alpha,\beta \in \Gamma$.
In particular, there exist $\alpha_0,\beta_0 \in \Gamma$ such that
$[A,\alpha_0,A,\beta_0,A] \in P$.
By primeness, at least one of the three entries in this interaction
must belong to $P$.
Since all three are equal to $A$, we conclude that $A \in P$.
Thus $P$ is semiprime.
\end{proof}

From a chemical perspective, this result indicates that in a prime
chemical ideal, any state whose self-interactions are entirely trapped 
within the ideal must itself be regarded as belonging to that ideal.
Prime subsystems therefore exclude the possibility of persistent 
external states whose internal dynamics is indistinguishable, in terms 
of reaction products, from that of genuine internal states.

The theory of prime and semiprime chemical ideals provides a way to 
decompose a TGS-chemical system into structurally meaningful components,
reflecting how reaction activity and mediated transformations are
distributed across the state space. Comparable decomposition principles appear in algebraic treatments of 
semirings and their ideal lattices \cite{HebischWeinert1998}.
.
In subsequent work, one may associate to a given system an appropriate
spectrum of prime chemical ideals and study its topology, thus
connecting the present framework with geometric methods.

\section{Homomorphisms of Chemical TGS}

Homomorphisms provide a natural mechanism for comparing different
TGS-chemical systems and transporting reaction behaviour from one system
to another.  
Just as homomorphisms of semirings or semigroups preserve algebraic
structure, homomorphisms of ternary~$\Gamma$-semirings preserve the
mediated ternary interaction that encodes chemical transformation.
The present section formalizes this notion and explains its chemical
significance.

Throughout, $(S,\Gamma,[\,])$ and $(S',\Gamma,[\,]')$ denote two
TGS-chemical systems sharing the same parameter set~$\Gamma$.
The requirement of a common $\Gamma$ reflects that mediators 
(catalysts, solvents, environmental conditions) are interpreted as 
parameters intrinsic to the interaction law and therefore must be 
preserved.

\subsection{Definition}

\begin{definition}
A map 
\[
  f \colon S \longrightarrow S'
\]
is called a \emph{homomorphism of TGS-chemical systems} if for all
$A,B,C \in S$ and all $\alpha,\beta \in \Gamma$,
\[
  f([A,\alpha,B,\beta,C])
  \;=\;
  [f(A),\alpha,f(B),\beta,f(C)]'.
\]
\end{definition}
This condition is analogous to structure-preserving maps in classical 
semiring and algebraic systems, where homomorphisms preserve the 
underlying interaction laws (see \cite{Golan1999,HebischWeinert1998}).

Thus, $f$ commutes with the ternary $\Gamma$-operation: applying the
reaction operation in~$S$ and then mapping the result via~$f$ yields the
same state as first mapping the inputs via~$f$ and then applying the
reaction operation in~$S'$.
In other words, $f$ is a structure-preserving transformation of chemical
environments.

Several immediate properties follow directly from the definition.

\begin{proposition}
Let $f \colon S \to S'$ be a TGS-homomorphism.
\begin{enumerate}[label=\textup{(\alph*)}]
  \item If $R \subseteq S$ is reaction-closed, then $f(R)$ is
        reaction-closed in~$S'$.
  \item If $I \subseteq S$ is a chemical ideal, then $f(I)$ is a 
        chemical ideal in~$S'$.
  \item If $J \subseteq S$ is a $\Gamma$-ideal of any type
        (left, right, middle, or two-sided), then $f(J)$ is a 
        $\Gamma$-ideal of the corresponding type in~$S'$.
\end{enumerate}
\end{proposition}

\begin{proof}
Each property is verified by direct substitution using the homomorphism 
identity.
For example, if $A,B,C \in R$ and $\alpha,\beta \in \Gamma$, then the
reaction-closedness of $R$ gives $[A,\alpha,B,\beta,C] \in R$, and
applying~$f$ yields
\[
  f([A,\alpha,B,\beta,C])
  =
  [f(A),\alpha,f(B),\beta,f(C)]' \in f(R),
\]
establishing reaction-closedness of $f(R)$.
The remaining cases follow the same pattern.
\end{proof}

This result shows that homomorphisms are compatible with the structural
subsystems developed in Section~4: reactors, basins, and pathways are
mapped to reactors, basins, and pathways in the target system.

\subsection{Chemical meaning}

A TGS-homomorphism models a consistency-preserving transformation 
between chemical environments.
Its chemical interpretations include the following:

\paragraph*{(1) Change of solvent or medium.}
Suppose $S$ describes reaction behaviour in solvent~$X$ and $S'$ in
solvent~$Y$.
A homomorphism 
\( f \colon S \to S' \)
represents a map translating chemical states from the $X$-environment to
the $Y$-environment such that the mediated interactions are preserved:
a triple interaction in $X$ corresponds exactly to the mapped triple
interaction in $Y$. This formalizes the intuitive idea that a well-defined solvent change
should send reaction pathways to reaction pathways without altering
their essential structure.

\paragraph*{(2) Change of catalyst or catalytic regime.}
Different catalytic environments can be modelled by different 
TGS-chemical systems built on the same parameter space~$\Gamma$ but with
distinct state spaces or distinct ternary interaction laws.
A homomorphism 
\[
  f \colon S \to S'
\]
can represent the adjustment of reaction behaviour when switching from
one catalyst to another.
The preservation of the mediated operation ensures that catalytic 
effects are transferred systematically rather than arbitrarily.Such environment-to-environment mappings have analogues in algebraic 
treatments of ternary and parameter-dependent transformations 
(see \cite{Tropashko2006}).

\paragraph*{(3) Controlled mapping between chemical environments.}
More generally, a homomorphism encodes any structured change of 
environment where reaction behaviour is transformed coherently.
This may represent, for example:
\begin{itemize}
  \item embedding a system with restricted state space into a larger one;
  \item coarse-graining a complex reaction network into a simpler model;
  \item mapping between different thermodynamic or field-controlled
        environments;
  \item abstraction from microscopic to effective macroscopic states.
\end{itemize}
In each of these examples, the homomorphism ensures that reaction 
mechanisms and mediator influences retain their form under translation.

\paragraph*{(4) Compatibility with reaction pathways.}
Since homomorphisms preserve the ternary $\Gamma$-operation, they also
preserve reaction pathways in the sense of Section~4.
Every reaction sequence 
\[
  X_0 \to X_1 \to \cdots \to X_n
\]
in~$S$ is carried by~$f$ to a reaction pathway
\[
  f(X_0) \to f(X_1) \to \cdots \to f(X_n)
\]
in~$S'$.
Thus, homomorphisms provide a bridge between dynamical behaviours in 
different systems, enabling the systematic study of how pathways 
transform under environmental changes.

\vspace{2mm}

Overall, homomorphisms of TGS-chemical systems play a role analogous to
structure-preserving maps in algebra, but their chemical interpretation
is richer: they express how reaction laws, mediators, and transformation
dynamics behave under coherent changes of environment.
This makes them powerful tools for both mathematical analysis and 
chemical modelling.

\section{Examples from Chemistry}

In this section we present several abstract but chemically meaningful
examples illustrating how mediated ternary interactions arise naturally
in chemical systems.In this section we present several abstract but chemically meaningful
examples illustrating how mediated ternary interactions arise naturally
in chemical systems.
These examples parallel mathematically formal approaches to chemical 
structure and transformation found in classical mathematical chemistry 
(see \cite{Balaban1992,Trinajstic1992,GutmanPolansky1986}).

The purpose of these examples is not to model specific experimental
systems but to show how familiar chemical phenomena can be expressed
within the TGS framework introduced above.

Throughout, $(S,\Gamma,[\,])$ denotes a TGS-chemical system in the sense
of Section~3, where $S$ represents chemical states and $\Gamma$ 
represents mediating conditions.

\subsection{Catalyzed reactions}

Catalysis provides a direct example of a mediated transformation in
which the presence of a catalyst modifies the reaction pathway without
being consumed.
Let $A,B,C \in S$ denote chemical states that participate in a
multi-step reaction, and let $\alpha,\beta \in \Gamma$ represent 
catalytic regimes.

Consider the ternary operation
\[
  [A,\alpha,B,\beta,C] = D.
\]
Here, $A$ may be interpreted as an initial reactant state, $B$ as an 
interacting partner or intermediate, and $C$ as a subsequent state 
through which the system passes.
The mediator $\alpha$ can encode the presence of a catalyst that opens
a specific reaction pathway, while $\beta$ may represent a co-catalyst 
or a secondary catalytic condition.

If $\alpha$ corresponds to a catalyst that lowers the effective barrier
between $A$ and $B$, and $\beta$ indexes a catalytic effect acting on
the transformation from $B$ to $C$, then $D$ represents the resulting
state of the catalyzed sequence.  
Different choices of $\alpha$ and~$\beta$ generally produce different 
outcomes:
\[
  [A,\alpha_1,B,\beta,C] \neq [A,\alpha_2,B,\beta,C],
\]
even when the underlying species $A,B,C$ remain fixed.
This expresses, in algebraic form, the well-known fact that changing 
catalysts can modify the reaction pathway or final products while 
preserving stoichiometry.

The ternary structure is essential here: the catalyst is not appended
externally but serves as an intrinsic argument of the reaction law.

\subsection{Phase transitions under thermodynamic control}

Phase transformations depend sensitively on thermodynamic parameters
such as temperature and pressure.
In the TGS framework, such environmental conditions are naturally
represented as elements of~$\Gamma$.

Let {$\Gamma$} consist of pairs $(T,p)$ corresponding to permissible
temperature--pressure regimes.Let $A,B,C \in S$ represent physical states of a substance, such as configurations or phase descriptors.
A ternary interaction
\[
  [A,(T_1,p_1),B,(T_2,p_2),C]
\]
produces a state $D$, where the mediators $(T_1,p_1)$ and $(T_2,p_2)$
govern the transitions between $A \to B$ and $B \to C$ respectively.

For example:
\begin{itemize}
  \item If $(T_1,p_1)$ represents conditions favouring melting,
        and $(T_2,p_2)$ represents conditions favouring vaporization,
        then $D$ may correspond to a higher-energy phase.
  \item If $(T_1,p_1)$ lies in a stability region for a solid phase,
        and $(T_2,p_2)$ lies in a stability region for a metastable 
        phase, then $D$ may encode a metastable state reached by 
        sequential transitions.
\end{itemize}

The ternary formulation captures the fact that multi-step phase 
transformations are governed not only by initial and final conditions
but also by intermediate thermodynamic regimes.
Different paths through $(T,p)$-space yield different outcomes, and the
dependence is faithfully recorded by the mediators in the 
$\Gamma$-operation.

\subsection{Quantum state transitions under external fields}

Quantum systems subject to external electromagnetic fields provide a 
further setting in which ternary, parameter-dependent interactions 
arise.
Let $S$ denote a set of quantum states, which may include electronic, 
vibrational, or spin configurations.
Let $\Gamma$ index external field parameters such as field strength,
frequency, or polarization.Such field-mediated transitions have abstract algebraic analogues in 
parameter-dependent ternary relation frameworks 
(see \cite{Tropashko2006}).

A ternary interaction
\[
  [A,\alpha,B,\beta,C] = D
\]
may then model a sequence of field-induced transitions:
\begin{itemize}
  \item $A \to B$ mediated by field parameter $\alpha$,
  \item $B \to C$ mediated by field parameter $\beta$,
  \item resulting in a state $D$ after the composite process.
\end{itemize}

For instance:
\begin{itemize}
  \item $\alpha$ may represent a low-frequency field inducing a 
        transition from $A$ to $B$;
  \item $\beta$ may represent a high-frequency field inducing a 
        transition from $B$ to $C$;
  \item the final state $D$ depends on the combined effect of both 
        fields in sequence.
\end{itemize}

The value of $D$ may differ significantly from what is obtained by 
either field alone, reflecting the well-established sensitivity of
quantum transitions to external field combinations.
The ternary structure captures this dependence by integrating the 
field parameters directly into the reaction law.

\vspace{2mm}

These examples illustrate how the ternary~$\Gamma$-operation provides a 
natural formalism for expressing catalysis, thermodynamic control, and 
field-induced quantum transitions within a single coherent algebraic
framework.
The examples are intentionally abstract, focusing on the structural
features that make TGS-chemical systems flexible enough to encode a wide
range of chemical behaviour.

\section{Conclusion and Future Work}

In this work we have developed an axiomatic framework for modelling
chemical systems using the structure of a ternary~$\Gamma$-semiring.
Beginning with the observation that chemical transformations are 
inherently multi-parameter and multi-state processes, we formulated a 
reaction law in which chemical states and mediating conditions appear as 
intrinsic arguments of a ternary operation.
This contrasts with the classical binary perspective, where catalysts 
and environmental factors are appended externally rather than 
participating structurally in the transformation process.

The foundational contribution of the paper lies in isolating the 
mathematical axioms that govern such mediated transformations and 
demonstrating how these axioms admit chemically meaningful 
interpretations.These axioms extend the structural principles familiar from classical 
semiring theory (see \cite{Golan1999,Bourne1951}) to a ternary 
$\Gamma$-mediated setting appropriate for chemical applications.

The ternary operation encodes multi-step transformations, the 
$\Gamma$-parameters incorporate catalytic and environmental effects, and 
the associativity and distributivity relations reflect coherence of 
reaction pathways.
The resulting concept of a TGS-chemical system provides a unified 
formalism in which multi-state, catalyst-dependent, and 
environment-dependent phenomena can be described algebraically.

We also developed the structural theory of these systems, introducing 
chemical ideals, $\Gamma$-ideals, and their prime and semiprime 
variants.
These notions identify chemically stable subsystems, reaction-closed 
domains, and regions whose internal behaviour governs the structure of 
mediated interactions.
Reaction pathways were characterized in terms of iterated ternary 
operations, and we showed how homomorphisms between TGS-chemical systems 
provide consistency-preserving maps between different chemical 
environments.
Finally, we illustrated the framework with abstract examples drawn from 
catalysis, phase transitions, and field-driven quantum processes.

\subsection*{Future directions}
The present framework opens several avenues for further development.

\begin{itemize}

  \item \textbf{Kinetic and dynamical refinements.}
  While the TGS formalism captures structural relationships between 
  states and mediators, incorporating explicit temporal or kinetic 
  data would allow the construction of mediated dynamical systems.
  A natural direction is to study sequences of ternary interactions as 
  discrete dynamical processes and to identify stability, periodicity, 
  or convergence phenomena within this setting.

  \item \textbf{Quantitative extensions.}
  The current theory treats $S$ and $\Gamma$ abstractly.
  Enriching these sets with additional algebraic or topological 
  structure—such as orders, metrics, or weights—could allow the 
  encoding of reaction energetics, field strengths, or graded 
  catalytic effects.
  Such extensions would be essential for connecting the TGS framework 
  to numerical models.

  \item \textbf{Categorical and geometric viewpoints.}
  The ideal theory developed here suggests the possibility of defining 
  spectra of prime chemical ideals and studying their geometric 
  features.
  This may lead to a form of \emph{ternary $\Gamma$-geometry} in which 
  chemical structure is represented through geometric invariants of the 
  spectrum.

  \item \textbf{Computational and AI-based models.}
  The unified treatment of states and mediators makes TGS-chemical 
  systems natural candidates for symbolic or rule-based computational 
  models.
  Subsequent work may explore how the ternary operation interacts with 
  algorithmic reasoning, abstraction, or machine-assisted simulation, 
  thereby linking algebraic chemistry with emerging methodologies in 
  computational chemistry and symbolic AI.Such directions resonate with semiring-based computational frameworks 
and ternary parameterized transformations explored in abstract algebraic 
settings (see \cite{Tropashko2006}).

\end{itemize}

\vspace{2mm}

Overall, the ternary~$\Gamma$-semiring viewpoint offers a flexible and 
conceptually coherent foundation for the algebraic study of chemical 
systems.
The theory developed in this paper establishes the basic structure on 
which further analytical, geometric, and computational developments can 
be built.

\end{document}